Exercise 12-8 Volume Trade-Off Decisions [LO12-5, LO12-6]

Barlow Company manufactures three products—A, B, and C. The selling price, variable costs, and contribution margin for one unit of each product follow:

Product

A B C
Selling price $ 160 $ 270 $ 210
Variable expenses:
Direct materials 16 80 24
Other variable expenses 108 90 144
Total variable expenses 124 170 168
Contribution margin $ 36 $ 100 $ 42
Contribution margin ratio 23 % 37 % 20 %

The same raw material is used in all three products. Barlow Company has only 5,700 pounds of raw material on hand and will not be able to obtain any more of it for several weeks due to a strike in its supplier’s plant. Management is trying to decide which product(s) to concentrate on next week in filling its backlog of orders. The material costs $8 per pound.

Required:

1. Calculate the contribution margin per pound of the constraining resource for each product.

2. Assuming that Barlow has unlimited demand for each of its three products, what is the maximum contribution margin the company can earn when using the 5,700 pounds of raw material on hand?

3. Assuming that Barlow’s estimated customer demand is 500 units per product line, what is the maximum contribution margin the company can earn when using the 5,700 pounds of raw material on hand?

4. A foreign supplier could furnish Barlow with additional stocks of the raw material at a substantial premium over the usual price. Assuming Barlow’s estimated customer demand is 500 units per product line and that the company has used its 5,700 pounds of raw material in an optimal fashion, what is the highest price Barlow Company should be willing to pay for an additional pound of materials?
Answer #1

1) Contribution margin per pound

A B C
Contribution margin per unit 36 100 42
Pound per unit 2 10 3
Contribution margin per pound 18 10 14

2) Total Contribution margin = 5700*18 = $102600

3) Contribution margin

Pound Unit Contribution margin
A 500*2 = 1000 500 1000*18 = 18000
C 500*3 = 1500 500 1500*14 = 21000
B 3200 320 3200*10 = 32000
Total 5700 71000

Total Contribution margin = $71000

4) Price per pound = 10+8 = 18 per pound